We got demand function q=6000-30p (q quantity, p=price). The total cost function is C(q)=72000+60q a) Solve the demand function for p and define its domain b) Find the marginal cost funtion MC and explain the result c) Find the income function R(q)
1
Expert's answer
2020-05-29T09:27:15-0400

a) demand function for p:

"P=200-Q\/30."

Domain "0<=Q<=6 000".

b)MC is a derivative of a total cost function, so

"MC=60."

c) "R = P*Q-C"

"R=(200-Q\/30)*Q-72000-60Q=140Q-Q^2\/30-72000."